OpenStudy (anonymous):

what is the slope of 6y=4(x+2)

OpenStudy (goformit100):

Warm Welcome To Open Study ......... Sir/ Ma'am @devera

OpenStudy (mathmale):

At this stage in your learning Algebra, the appropriate approach to finding the slope of 6y=4(x+2) is to solve this equation for y, and then to re-write the result in the form y=mx+b. Would you try this, please? What is the coefficient of x?

OpenStudy (anonymous):

6y=4(x+2) 6y=4x+8 so 4 is the slope

OpenStudy (mathmale):

@katlee15: Thank you for trying to help. Note that the preferred way of helping others on OpenStudy is to guide them through the solution of their own problems by themselves. In other words, we don't do the work of others for them. secondly, 4 is not the correct slope here. Try doing this problem over again, in private (not here). @devera: Would you please respond to my previous post? It was addressed to you.